OpenStudy (anonymous):

Find the number b such that the line y = b divides the region bounded by the curves y = x^2 and y = 4 into two regions with equal area. Give your answer correct to 3 decimal places.

OpenStudy (agl202):

9x² = 25 .......... find the intersections x = ± 5/3 A = 2 ∫ 25 - 9x² dx [0,5/3] ........ area bounded by y = 25 and y = 9x² = 500/9 9x² = b .......... find the intersections x = √b/3 250/9 = 2 ∫ b - 9x² dx [0,√b/3] .... half the area 250/9 = 2bx - 6x^3 [0,√b/3] 250/9 = 4/9 b^(3/2) b = 25/∛4 Answer: b = 25/∛4 .... or approx. 15.74901312 Hope this helps u!!! =)

OpenStudy (anonymous):

thank u!!!!!!!!!!!!!! :)

OpenStudy (agl202):

Anytime!

OpenStudy (anonymous):

can u plz help me with another one? :)